# exa-AMD: An Exascale-Ready Framework for Accelerating the Discovery and Design of Functional Materials

### v1

@@ 4 Performance and Scalability
The exa-AMD framework is explicitly designed to leverage modern high-performance computing (HPC) facilities, including exascale clusters with both CPU and GPU architectures with the use of Parsl. By decoupling workflow logic from execution resources, exa-AMD can efficiently multiplex tens of thousands of independent jobs, automatically scaling workload to available nodes and managing asynchronous dependencies. Importantly, Parsl supports dynamic provisioning and workload elasticity: as the workflow proceeds through different stages in the workflow, computing resources can be grown or shrunk, and jobs can transparently recover or resume from failures without user’s intervention. This ensures efficient backfilling and utilization of large, shared supercomputers.
@@
To quantify performance, we conducted systematic scaling tests three representative systems: Na-B-C, Ce-Co-B, and Fe-Co-Zr. Na-B-C only involves light elements, making it the simplest system among them. Ce-Co-B is relatively complicated, involving rare-earth element as well as 3d transition metal. Fe-Co-Zr is selected as a typical rare-earth-free magnetic system. For each system, we measured total wall-clock time required to complete a workflow consisting of structure generation, CGCNN-based screening, and a fixed-size pool of parallelized DFT relaxations. Benchmarks were performed on both CPU and GPU partitions, with node counts up to 32 for Na-B-C system, and up to 256 nodes for Ce-Co-B and Fe-Co-Zr systems. [explain why 32 nodes for Na-B-C system with light element, and 256 nodes for complex systems]. Figs. 2, 3, and 4 summarize the strong scaling results: wall-clock times decrease nearly ideally with node count, following a 1/$N$ scaling, on both CPU and GPU backends, illustrating excellent parallel efficiency. We have performed the benchmark tests on two large computers, perlmutter from NERSC and Chicoma from Los Alamos National Laboratory. [details on GPU nodes, CPU nodes]. For example, in the Na-B-C system, wall-clock time reduced from 1550 minutes on a single node with 4 GPUs to 88 minutes (32 GPU nodes), and from 1520 to 98 minutes over the same range on CPUs. GPU benchmarks consistently achieve slightly faster runtimes than CPU. Moreover, compared to CPU benchmarks, GPU benchmarks exhibit strong scaling that is closer to ideal behavior at higher node counts, reflecting the efficiency of accelerating ML inference and ab initio calculations by modern GPU architectures.
Performance comparisons for the Ce-Co-B and Fe-Co-Zr systems reveal similar strong scaling: for Ce-Co-B, total workflow time dropped from 2112 to 50 minutes on GPUs (4 to 256 nodes), and 2091 to 118 minutes on CPUs (4 to 128 nodes) . For the more demanding Fe-Co-Zr system, wall-clock time decreased from 2768 to 63 minutes as GPU node counts were increased from 4 to 256, and from 3018 to 155 minutes as CPU node counts change from 4 to 128. The parallel efficiency typically remaining at above 80% across the tested range. The highly modular design managed by Parsl ensures that each job (whether a ML inference, structure relaxation, or post-processing stage) is distributed, tracked, and aggregated without significant idle time or manual intervention. Dynamic resource allocation and robust fault tolerance further maximize throughput and utilization, particularly in shared supercomputing environments with variable queueing latency and hardware availability.
@@
Figure 2: Strong scaling of exa-AMD workflow for the Na-B-C system. Wall-clock times are shown for both GPU and CPU nodes (1, 2, 4, 8, 16, 32). The workflow exhibits near-linear speed-up and substantial GPU acceleration at all node counts.
@@
Figure 3: Benchmark results for the Ce-Co-B system executed on GPU architectures for 4 to 256 nodes, and on CPU for 4 to 128 nodes. Performance demonstrates efficient strong scaling and consistent GPU advantage for large-scale campaigns.
@@
Figure 4: Wall-clock time as a function of node count for the Fe-Co-Zr system, on GPU architectures for 4 to 256 nodes, and on CPU for 4 to 128 nodes. . Strong scaling is maintained across the full range, underscoring exa-AMD’s readiness for large, exascale discovery campaigns.
@@
These results directly demonstrate that exa-AMD achieves almost optimal scaling and efficient throughput across diverse computational environments. This enables million-structure ML screening in under an hour and ab initio relaxation of thousands of candidates within a day on modern clusters.
To provide deeper insight into exa-AMD’s resource utilization, we further analyzed the time distribution among major phases in the workflow using Na-B-C as an example. As illustrated in Fig. 5, the vast majority of the total wall-clock time is consumed by VASP calculations (97.6 %). Structure generation and selection, the CGCNN machine learning inference, and selection/filtering steps together require under 3% combined.
@@
Figure 5: Time distribution among major workflow phases for the Na-B-C benchmark. Shown are the fractions for structure generation (blue), CGCNN inference (orange), structure selection (green), and VASP DFT calculations (red). Nearly all computational effort is spent on the high-throughput ab initio tasks, underscoring the efficiency of accelerated machine learning screening and workflow management stages.
@@
With GPU acceleration, the CGCNN model can process over 1 million structures within minutes, allowing exa-AMD to rapidly down-select candidates and prioritize compute resources for only the most promising compounds. This efficient distribution—made possible by Parsl’s fine-grained, parallel task scheduling—demonstrates that the workflow overhead is minimal and that exa-AMD achieves near-optimal throughput for production-scale discovery, limited primarily by the computationally intensive quantum mechanical calculations.
This workflow profiling confirms that by using advanced job control and asynchronous task management, non-ab initio workflow steps incur negligible bottlenecks, even as campaign size and node count increase. Resources are automatically released from structure screening phases and reallocated for the surge in DFT job submissions, allowing seamless scaling and minimized overall turnaround time for large discovery efforts.
@@ 5 Example Applications
A compelling demonstration of the exa-AMD framework is its application to accelerated discovery and design of rare-earth-free permanent magnets in the ternary Fe-Co-Zr system. One of the hallmarks of exa-AMD is its user-focused interface and workflow: the user simply specifies a set of chemical elements—in this case, Fe, Co, and Zr—and the directory containing the prototype structures, and the framework automates the entire process of generating, screening, and validating hypothetical compounds, requiring minimal manual intervention.
@@
After specifying the elements, the workflow proceeds by generating a broad pool of candidate structures using database-driven template substitution, combinatorial permutation, and lattice scaling. CGCNN predicts the formation energies for nearly 900,000 candidate Fe-Co-Zr compositions in under 15 minutes using a single GPU node. Structures with predicted formation energy below a set threshold (e.g., $E_{\mathrm{f}}<-0.1$ eV/atom) are automatically filtered in. After removing duplicated structures, this yields about 3,100 distinct candidates for ab initio validation. The subsequent first-principles (VASP) relaxation of all 3,100 structures was completed in less than 12 hours on a modest GPU cluster. The only user input required throughout this discovery pipeline was the choice of target elements. All stability screening, filtering, relaxation, and post-processing were handled automatically, demonstrating both the simplicity and power of the exa-AMD approach.
@@
As a final output, exa-AMD provided significantly updated convex hulls for the Fe-Co-Zr phase diagram (Fig. 6), as well as an annotated list of newly predicted stable and low-energy (metastable) ternary compounds. Nine new stable Fe-Co-Zr phases and 81 promising metastable compounds (within 0.1 eV/atom above the hull) were discovered, dramatically extending the known landscape for this key system. For selected candidates, exa-AMD also provided DFT-optimized crystal structures, site-specific magnetic properties, and both electronic band structures and densities of states (as in Table 1 and supplementary figures of the Fe-Co-Zr study).
Figure 6: The updated convex hull with the newly predicted Fe-Co-Zr compounds by DFT calculations based on the structures selected through exa-AMD framework. Red dots are the 9 newly predicted stable phases; colored triangles are the 81 newly identified unique metastable phases with low formation energies (within 0.1 eV/atom above the updated convex hull).
The exa-AMD workflow is equally applicable to a wide range of materials and other target properties. For demonstration, in the Na-B-C system, the user again simply specifies the three desired elements. exa-AMD then systematically explores all possibilities generated by template substitution, screens stable candidates by ML, and refines selected compound energies and properties by DFT calculations. The outputs include both updated convex hulls (for phase stability visualization) and, if desired, relaxed structure information for experimental synthesis or further theoretical study. Fig. 7 shows a sample output for Na-B-C. The discovery of quaternary systems and more complex systems can be pursued using the same high-throughput approach.
@@
Figure 7: The updated convex hull with the newly predicted Na-B-C compounds by DFT calculations based on the structures selected through exa-AMD framework. Colored squares are identified metastable phases with low formation energies (within 0.2 eV/atom above the convex hull).
This modular, scalable, and element-agnostic strategy allows exa-AMD users to efficiently perform reliable high-throughput discovery campaigns for virtually any combination of elements, fulfilling a critical need for rapid, reproducible computational materials design.
@@ 6.1 Installation
  • 1.
    Clone the repository from https://github.com/ml-AMD/exa-amd/ and follow the conda installation instructions for exa-AMD and its dependencies (conda env create -f amd_env.yml).
  • 2.
    Prepare your computational environment with compatible quantum simulation packages (e.g., VASP, Quantum ESPRESSO).

### v2

@@ 4 Performance and Scalability
With the use of Parsl, the exa-AMD framework is explicitly designed to leverage modern high-performance computing (HPC) facilities, including exascale or heterogeneous clusters with both CPU and GPU architectures. By decoupling workflow logic from execution resources, exa-AMD can efficiently multiplex tens of thousands of independent jobs, automatically scaling workload to available nodes and managing asynchronous dependencies. Importantly, Parsl supports dynamic provisioning and workload elasticity: as the workflow proceeds through different stages in the workflow, computing resources can be grown or shrunk, and jobs can transparently recover or resume from failures without user’s intervention. This ensures efficient backfilling and utilization of large, shared supercomputers.
@@
To quantify performance, we conducted systematic scaling tests on three representative systems: Na-B-C, Ce-Co-B, and Fe-Co-Zr. Na-B-C only involves light elements, making it the simplest system among them. Ce-Co-B is relatively complicated, involving rare-earth element as well as 3d transition metal. Fe-Co-Zr is selected as a typical rare-earth-free magnetic system. For each system, we measured total wall-clock time required to complete a workflow consisting of structure generation, CGCNN-based screening, structure selection and a fixed-sized pool of parallelized DFT relaxations. Structure generation and selection were performed on CPUs only, while CGCNN screening and DFT calculations made use of both CPU and GPU architectures depending on the benchmark configuration. Benchmarks were performed on both CPU and GPU partitions, with node counts up to 32 for the Na-B-C system, and up to 256 for the Ce-Co-B and Fe-Co-Zr systems. The significantly smaller node count for Na-B-C reflects the lower computational demands of this system: it involves only light elements (Na, B, C) with fewer electrons and simpler electronic structures, resulting in faster DFT convergence and shorter computation time per structure. In contrast, Ce-Co-B and Fe-Co-Zr are substantially more computationally expensive—Ce-Co-B contains rare-earth element Ce with partially filled 4f orbitals requiring careful electronic treatment, while Fe-Co-Zr involves 3d transition metals with complex magnetic interactions. These systems require more self-consistent-field iterations for convergence, necessitating larger node counts to achieve comparable wall-clock times for benchmarking purposes.
Figs. 4, 5, and 6 summarize the strong scaling results: wall-clock times decrease nearly ideally with node count, following a $1/N$ scaling, on both CPU and GPU backends, illustrating excellent parallel efficiency. We performed the benchmark tests on two large computers, Perlmutter at the National Energy Research Scientific Computing Center (NERSC) and Chicoma at Los Alamos National Laboratory (LANL). Each Perlmutter’s CPU node is comprised of two AMD EPYC 7763 (Milan) 64-core 2.45 GHz CPUs; each GPU node has one CPU of the same architecture and four NVIDIA A100 GPUs. Chicoma has similar hardware specifications: each CPU node has two AMD EPYC 7H12 (Rome) 64-core 2.6 GHz CPUs; whereas each GPU node has one AMD EPYC 7713 (Rome) 64-core 2 GHz CPU and four NVIDIA A100 GPUs. For example, in the Na-B-C system on NERSC’s Perlmutter supercomputer, wall-clock time reduced from 1550 minutes on a single node with 4 GPUs to 88 minutes on 32 GPU nodes, and from 1520 to 98 minutes over the same number of CPU nodes. GPU benchmarks consistently achieve slightly faster runtime than on CPUs. Moreover, compared to CPU benchmarks, GPU benchmarks exhibit strong scaling that is closer to ideal behavior at higher node counts, reflecting the efficiency of accelerating ML inference and ab initio calculations by modern GPU architectures.
@@
Figure 4: Strong scaling of exa-AMD workflow for the Na-B-C system on NERSC’s Perlmutter supercomputer. Wall-clock times are shown for both GPU and CPU nodes (1, 2, 4, 8, 16, 32). The workflow exhibits near-linear speed-up and substantial GPU acceleration at all node counts.
@@
Figure 5: Benchmark results for the Ce-Co-B system on NERSC’s Perlmutter supercomputer, executed on GPU architectures for 4 to 256 nodes, and on CPU for 4 to 128 nodes. Performance demonstrates efficient strong scaling and consistent GPU advantage for large-scale campaigns.
@@
Figure 6: Wall-clock time as a function of node count for the Fe-Co-Zr system on NERSC’s Perlmutter supercomputer, on GPU architectures for 4 to 256 nodes, and on CPU for 4 to 128 nodes. Strong scaling is maintained across the full range, underscoring exa-AMD’s readiness for large, exascale discovery campaigns.
Performance comparisons for the Ce-Co-B and Fe-Co-Zr systems on NERSC’s Perlmutter supercomputer reveal similar strong scaling: for Ce-Co-B, total workflow time dropped from 2112 to 50 minutes on GPUs (4 to 256 nodes), and 2091 to 118 minutes on CPUs (4 to 128 nodes). For Ce-Co-B and Fe-Co-Zr benchmarks, we start from 4 nodes rather than the conventional single-node baseline due to practical time constraints. A single-node run would require approximately 2.5–3 days to complete the full workflow for these computationally demanding systems. Such extended runtime are impractical for systematic benchmarking studies and exceed typical time limits on shared HPC resources. By using 4 nodes as our baseline, we maintain reasonable benchmark completion times while still demonstrating strong scaling behavior across two orders of magnitude in node count. The near-ideal $1/N$ scaling observed from 4 to 256 nodes strongly suggests that this scaling behavior would extend to lower node counts, though at the cost of prohibitively long wall-clock times for practical use. The parallel efficiency typically remains at above 80% across the tested range. The highly modular design managed by Parsl ensures that each job (whether a ML inference, structure relaxation, or post-processing stage) is distributed, tracked, and aggregated without significant idle time or manual intervention. Dynamic resource allocation and robust fault tolerance further maximize throughput and utilization, particularly in shared supercomputing environments with variable queueing latency and hardware availability.
@@
These results directly demonstrate that exa-AMD achieves almost optimal scaling and efficient throughput across the CPU and GPU machines we tested on. This enables million-structure ML screening in under an hour and ab initio relaxation of thousands of candidates within a day on modern clusters.
To examine the time distribution of different phases of the exaAMD workflow, we analyzed the Na-B-C system as a representative case (Fig. 7). First principles calculations performed with VASP dominate the total execution time, representing about 90.5% on 16 nodes. The remaining time is divided among structure generation, CGCNN inference for formation energy prediction, and structure selection. The structure selection phase is parallelized only with shared memory and therefore runs on a single node, which limits its scalability relative to the other stages. Nevertheless, even at large scales this phase remains a minor contributor to the total execution time, confirming that the overall cost is governed primarily by the ab initio calculations.
(a) Proportion of time spent on DFT in an end-to-end workflow.
(b) Pre-DFT timing breakdown (absolute time).
@@
Figure 7: Time distribution among major workflow phases for the Na-B-C benchmark performed on LANL’s Chicoma supercomputer. Shown are the fractions for structure generation (blue), CGCNN inference (orange), structure selection (green), and VASP DFT calculations (red). For clarity, we show the exact time for pre-DFT calculations, and only the relative percentage (compared to total execution time) for DFT calculations.
@@
With GPU acceleration, the CGCNN model can process over 1 million structures within minutes, allowing exa-AMD to rapidly down-select candidates and prioritize compute resources for only the most promising compounds. This efficient scalability—made possible by Parsl’s fine-grained, parallel task scheduling—demonstrates that the workflow overhead is minimal and that exa-AMD achieves near-optimal throughput for production-scale discovery, limited primarily by the computationally intensive quantum mechanical calculations.
As the workflow scales to larger node counts, the relative overhead from non-DFT tasks (structure generation, ML inference, and job management) increases modestly due to communication and coordination costs, reaching approximately 15% at 32 nodes for the Na-B-C system. However, this scaling behavior remains acceptable for production use: the absolute time spent on these tasks decreases with parallelization, and the overall workflow still exhibits strong scaling efficiency (¿80%) across the tested range. By leveraging Parsl’s asynchronous task management and dynamic resource allocation, exa-AMD minimizes idle compute time and ensures that the bulk of computational resources remain focused on the high-value DFT calculations throughout the discovery campaign. Resources are automatically released from structure screening phases and reallocated for the surge in DFT job submissions, allowing seamless scaling and minimized overall turnaround time for large discovery efforts.
@@ 5 Example Applications
A compelling demonstration of the exa-AMD framework is its application to accelerated discovery and design of rare-earth-free permanent magnets in the ternary Fe-Co-Zr system [FeCoZr]. One of the hallmarks of exa-AMD is its user-focused interface and workflow: the user simply specifies a set of chemical elements—in this case, Fe, Co, and Zr—and the framework automates the entire process of generating, screening, and validating hypothetical compounds using the prototype structures provided by exa-AMD by default, requiring minimal manual intervention.
@@
After specifying the elements, the workflow proceeds by generating a broad pool of candidate structures using elemental substitution on template structures and lattice scaling. CGCNN predicts the formation energies for nearly 900,000 candidate Fe-Co-Zr compositions in under 15 minutes using a single GPU node. Structures with predicted formation energy below a set threshold (e.g., $E_{\mathrm{f}}<-0.1$ eV/atom) are automatically filtered in. After removing duplicated structures, this yields about 3,100 distinct candidates for ab initio validation. Subsequent first-principles calculations using VASP, including structural relaxation and total energy calculations for all 3,100 structures, were completed in less than 12 hours on a single GPU node with 4 GPUs. The only user input required throughout this discovery pipeline was the choice of target elements. All stability screening, filtering, relaxation, and post-processing were handled automatically, demonstrating both the simplicity and power of the exa-AMD approach.
@@
As a final output, the exa-AMD framework provided a significantly updated convex hull for the Fe-Co-Zr phase diagram (Fig. 8). The calculated results revealed a dramatically extended known landscape for this key system, including the discovery of nine new stable Fe-Co-Zr phases and 81 promising metastable compounds (those within $0.1$ eV/atom above the hull). Beyond phase stability predictions, exa-AMD provides DFT-optimized crystal structures for these new compounds. Optional postprocessing can be employed for selected candidates to calculate complex materials properties, such as site-specific magnetic properties, electronic band structures, and densities of states when required [FeCoZr].
(a) Known Fe-Co-Zr convex hull.
(b) Updated Fe-Co-Zr convex hull.
@@
Figure 8: Comparison of the Fe-Co-Zr convex hull. (a) The convex hull calculated using previously known stable structures from the Materials Project [MP]. (b) The updated convex hull from the exa-AMD framework, showing newly predicted stable and metastable phases. Colored squares in (b) indicate newly identified metastable phases with low formation energies (within 0.1 eV/atom above the convex hull).
The exa-AMD workflow is equally applicable to a wide range of materials and other target properties. For demonstration, the Na-B-C system provides a clear example of the framework’s power in a complex chemical space (Fig. 9). The user again simply specifies the three desired elements. exa-AMD then systematically explores all possibilities generated by template substitution, screens stable candidates using ML models, and refines selected compound energies and properties by DFT calculations. The outputs include both updated convex hulls (for phase stability visualization) and, if desired, relaxed structure information for experimental synthesis or further theoretical study. exa-AMD revised the known thermodynamic landscape for this system, and discovered 11 new metastable Na-B-C phases (within the $0.1$ eV/atom threshold above the hull). The ability to efficiently sample and optimize structures across this ternary space allows us to pinpoint compounds that may possess desirable properties for applications such as energy storage or high-hardness materials.
(a) Known Na-B-C convex hull.
(b) Updated Na-B-C convex hull.
Figure 9: Comparison of the Na-B-C convex hull. (a) The previously established phase stability landscape calculated using structures obtained from the Materials Project [MP]. (b) The updated convex hull resulting from the exa-AMD framework. Colored squares in (b) indicate newly identified metastable phases with low formation energies (within 0.1 eV/atom above the convex hull).
The discovery of quaternary systems and higher-order multi-principal element alloys (e.g., quinary Al-Co-Cr-Fe-Ni high-entropy alloys) can be pursued using the same high-throughput approach. The framework is particularly valuable for exploring computationally demanding systems that challenge conventional DFT workflows: rare-earth-containing compounds (e.g., Ce-Co-B, Nd-Fe-B, Sm-Co alloys) with strongly correlated 4f electrons requiring careful convergence; actinide-based materials (e.g., U-Mo-Zr nuclear fuels) with 5f orbital complexity and large spin-orbit coupling; systems with heavy transition metals (e.g., Pt-Ir-Rh catalysts, W-Re-Os refractory alloys) demanding dense k-point meshes and high plane-wave cutoffs; and magnetic materials with complex spin configurations (e.g., frustrated magnets, non-collinear magnetic structures in Mn-based alloys). These more complex systems, which involve combinatorially larger composition spaces and substantially increased computational cost per structure, benefit significantly from exa-AMD’s automated workflow, GPU acceleration, and efficient scaling to hundreds of nodes, making them particularly suitable for exascale discovery campaigns. The modular, scalable, and element-agnostic strategy allows exa-AMD users to efficiently perform reliable high-throughput discovery campaigns for virtually any combination of elements, fulfilling a critical need for rapid, reproducible computational materials design.
@@ 6 User Guide
This section is not intended to replace the official online documentation of exa-AMD. Instead, it is meant to complement it by providing a concise overview of the installation, configuration, and execution steps, together with practical information that helps new users quickly identify and resolve common issues encountered when setting up or running the workflow.
@@ 6.1 Installation
exa-AMD is publicly available at https://github.com/ML-AMD/exa-amd. The package can be installed either from a prebuilt release or directly from source. The recommended method is to install the distributed wheel:
# create & activate a fresh virtual environment
# python -m venv myenv
# source myenv/bin/activate
# python -m pip install --upgrade pip
pip install "https://github.com/ML-AMD/exa-amd/releases/download/v0.1.1/exa_amd-0.1.1-py3-none-any.whl"
exa_amd --help
For development or testing, one can clone the repository, create the Conda environment defined in amd_env.yml, and execute the main module directly:
git clone https://github.com/ML-AMD/exa-amd.git
cd exa-amd
conda env create -f amd_env.yml
conda activate amd_env
export PYTHONPATH=\$(pwd):\$PYTHONPATH
export PYTHONPATH="\$(pwd)/ml_models:\$PYTHONPATH"
python exa_amd.py --help
The last command lists all the command line arguments that can be passed into the main executable to control the workflow setting. All software dependencies are specified in amd_env.yml at the root of the repository and are not repeated here. In particular, a working installation of a first-principles code such as VASP is required for the DFT optimization stages, together with the corresponding VASP input files, e.g., the projector-augmented-wave (PAW) pseudopotentials.
6.2 Configuration and Execution
User input is provided through a single JSON configuration file that defines the target chemical system, the relevant working directories, and a reference to a Parsl configuration file that specifies the computational resources. These parameters in the JSON file can be overridden by the command line arguments provided to the main executable exa_amd.py. For brevity, we do not reproduce the entire schema here since a complete description is available in the online documentation. Instead, we summarize below the most relevant parameters required for a new user to immediately begin using exa-AMD. As a starting point, users are encouraged to copy and adapt the configuration template configs/perlmutter.json distributed with the repository (this holds even if the target platform is not Perlmutter).
"work_dir" Absolute path to the working directory for the first three stages (structure generation, formation-energy prediction, and structure selection); typically on a scratch filesystem.
"elements" Target chemical composition: ternary or quaternary systems (e.g., "Na-B-C" or "Na-B-H-C").
"initial_ structures_dir" Directory of initial structure templates used to generate hypothetical compounds. A curated set is available at https://zenodo.org/records/17180192; users may also provide their own.
"parsl_config" Selects a named runtime resource configuration (e.g., "perlmutter_premium").
"parsl_configs_dir" Directory searched at startup for user-authored Parsl configuration files.
"vasp_std_exe" Executable name or absolute path of the VASP standard binary.
"vasp_work_dir" Working directory for VASP jobs; separate from "work_dir" to accommodate systems where scratch has constraints.
"vasp_pot_dir" Path to the PAW potentials directory for GGA calculations.
"vasp_output_file" Name of the CSV summarizing structures that successfully converged (e.g., "vasp_results.csv").
"vasp_nstructures" Maximum number of structures processed in this run (-1 = all). exa-AMD tracks progress and continues from the next structure on subsequent runs (e.g., 1–10, then 11–20).
"post_processing_ output_dir" Directory for post-processing outputs and summaries; if unset, post-processing is skipped.
"mp_rester_api_key" Materials Project API key for reference-data queries during post-processing (required if "post_processing_output_dir" is set).
The JSON file also defines general workflow parameters, such as the formation-energy threshold, worker counts, and VASP runtime options, and when using the built-in Parsl configurations it provides helper fields to set Parsl resources (for example, the number of nodes for the DFT calculations stage).
The Parsl settings above only select a configuration by name and indicate where configurations are discovered. We now outline how these configurations specify resource selection and executor behavior in practice. Indeed, a Parsl configuration describes how different workflow stages are mapped to CPU and GPU resources and how far each stage is allowed to scale. Below we provide a minimal illustrative snippet for a GPU-enabled VASP executor on Perlmutter from the PerlmutterConfig class, which also defines the different executors for other parts of the workflow (CGCNN prediction, structure generation, etc). The full configuration and complete examples are available in the Github repository.
class PerlmutterConfig(Config):
def __init__(self, json_config):
# VASP executor
vasp_executor = HighThroughputExecutor(
label=VASP_EXECUTOR_LABEL,
cores_per_worker=1,
available_accelerators=4,
provider=SlurmProvider(
account=gpu_account,
qos="premium",
constraint="gpu",
init_blocks=0,
min_blocks=0,
max_blocks=16,
nodes_per_block=1,
launcher=SimpleLauncher(),
walltime=’16:00:00’,
worker_init="moduleloadvasp/6.4.3-gpu",
In this example, we specify SlurmProvider to indicate that the computations will be executed on a Slurm-managed system. Parsl will create a Slurm script for job submission and requesting for computational resources. The fields account, qos, and constraint map directly to their srun/sbatch equivalents on Perlmutter. The setting max_blocks = 16 permits using up to sixteen GPU nodes (subject to the number of pending tasks), and available_accelerators = 4 schedules one task per GPU, for four concurrent VASP tasks per node. The label VASP_EXECUTOR_LABEL identifies the executor used for the DFT stage of the workflow, meaning that the relaxation and total-energy calculations are executed using the computational resources defined in this block. The other executors, used for structure generation, CGCNN inference, selection, and post-processing follow analogous definitions and are documented at https://ml-amd.github.io/exa-amd/workflow.html.
After defining a full configuration (see the parsl_configs directory at the root of the exa-AMD repository for complete examples), it must be registered so that it can be discovered at runtime:
register_parsl_config("perlmutter_premium", PerlmutterConfig)
The file that contains the registrations is then moved into a directory whose path is provided via the JSON parameter "parsl_configs_dir". This directory may contain multiple files, each with multiple configurations and calls to register_parsl_config; the only requirement is that registration names are unique across all files in the directory. At startup, exa-AMD reads all files in "parsl_configs_dir" and collects the registered configurations. The user selects one of them at runtime by setting the JSON parameter "parsl_config" (e.g., "perlmutter_premium").